# Data-dependent and Oracle Bounds on Forgetting in Continual Learning

### v2

@@ Abstract
In continual learning, knowledge must be preserved and re-used between tasks, maintaining good transfer to future tasks and minimizing forgetting of previously learned ones. While several practical algorithms have been devised for this setting, there have been few theoretical works aiming to quantify and bound the degree of Forgetting in general settings. We provide both data-dependent and oracle upper bounds that apply regardless of model and algorithm choice, as well as bounds for Gibbs posteriors. We derive an algorithm based on our bounds and demonstrate empirically that our approach yields tight bounds on forgetting for several continual learning problems.
@@ 1 Introduction
Continual learning is a burgeoning machine learning setting where data from different tasks are presented sequentially to the learner. The usual stated goal of methods in this setting is to adapt the learner to new tasks as they appear while also preserving its performance on previous tasks (De Lange et al., 2021; Hadsell et al., 2020; Parisi et al., 2019). This performance on previous tasks is called backward transfer, or forgetting, and one of the key challenges in continual learning is avoiding Catastrophic Forgetting (Goodfellow et al., 2014; Ramasesh et al., 2020; Kirkpatrick et al., 2017), meaning that performance on previous tasks degrades significantly as the model adapts to new tasks.
@@
Although avoiding catastrophic forgetting is desirable, much of the focus of continual learning research in recent years was on settings with a shared optimal solution. Realistically, we should not expect a single algorithm to perform optimally across all settings, for example if the data distribution changes gradually. A recent paper by Kumar et al. (2023) discusses continual learning as a computationally constrained optimization problem and argues that forgetting non-recurring information is not “catastrophic”, especially given changing environments. We will discuss this topic further in our empirical evaluation.
@@
While there have been several empirical approaches in the field, there are relatively few theoretical works that explore and attempt to quantify and bound this backward transfer. Some, such as Evron et al. (2022); Lin et al. (2023); Li et al. (2024) focus on linear models to consider the effect of task order and similarity on forgetting. Others, such as Bennani and Sugiyama (2020); Doan et al. (2021) utilize the NTK regime to focus on more complex task similarity measures as predictors of forgetting. Several more general works such as Benavides-Prado and Riddle (2022) apply notions of VC-dimension to arrive at more general scaling laws and upper bounds on forgetting for the ERM model, but may be difficult to apply for complex predictors due to the potentially large VC-dimension of models such as deep neural networks (Bartlett et al., 2019). We note that many of the known results (Bennani and Sugiyama, 2020; Evron et al., 2022; Benavides-Prado and Riddle, 2022) provide upper bound on forgetting for the training data. Our work, however, focuses on general bounds on forgetting for test data. We note that there are some previous results on forgetting for test data (Doan et al., 2021; Lin et al., 2023; Li et al., 2024) for specific model. In particular, Lin et al. (2023) provide an exact characterization of forgetting under specific assumptions in the linear regression setting.
@@
In this work, we explore upper bounds on forgetting that apply for both general and specific models. We use the PAC-Bayes (McAllester, 1999; Catoni, 2004; Alquier, 2024) framework to derive and analyze upper bounds on backward transfer, focusing on the Gibbs posterior (Casella and George, 1992). In Section 3 we derive general bounds for the two-task and $T$-task settings without making any model assumptions. We then focus our discussion in Section 4 on oracle bounds for the Gibbs posterior in general, and under specific assumptions of task similarity. We then proceed to derive an algorithm based on our bounds and we empirically validate it [^1] Anonymized Code is available in a separate zip file. on several continual learning problems in Section 5. To the best of our knowledge, our bounds are the first general PAC-Bayes upper bounds on forgetting in continual learning that are applicable regardless of model choice and task distribution.
@@ 2 Problem definition
The expected loss of a given hypothesis $h\in\mathcal{H}$ is defined as $\mathcal{L}(h,\mathcal{D})\triangleq E_{z\in\mathcal{D}}\ell(h,z)$. The empirical loss of a hypothesis w.r.t. a sample $S$ is defined as ^$\hat{\mathcal{L}}(h,S)\triangleq\frac{1}{m}\sum_{j=1}^{m}\ell(h,z_{j})$.
In the following Section, and in Section 4, we first consider only two tasks from distributions $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$, referring to source and target. Within the classic PAC-Bayes setting (McAllester, 1999; Alquier, 2024), let $Q_{s}$ be a distribution over the set of hypotheses learned by some algorithm $J_{s}$ over $S_{s}\sim\mathcal{D}_{s}$ and a data-free prior hypothesis distribution $P$, such that $Q_{s}=J_{s}(S_{s},P).$ We then proceed to utilize another algorithm $J_{t}$ that operates on $S_{t}\sim\mathcal{D}_{t}$, making use of $Q_{s}$, such that $Q_{s:t}=J_{t}(S_{t},Q_{s}).$ For now, we make no assumptions on $J_{s}$ and $S_{t}$ other than their inputs. Of particular note is that $J_{t}$ has access to information on the previous task only via the prior distribution $Q_{s}$.
@@
$F(Q_{s:t},\mathcal{D}_{s})$ $\triangleq\mathrm{BWT}(Q_{s:t},\mathcal{D}_{s})-\mathbb{E}_{h\sim Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})\right]$
$=\mathcal{L}(Q_{s:t},\mathcal{D}_{s})-\mathcal{L}(Q_{s},\mathcal{D}_{s}).$
Intuitively, the backward transfer $\mathcal{L}(Q_{s:t},\mathcal{D}_{s})$ measures the performance of the updated model on the previous task $\mathcal{D}_{s}$, and the forgetting $F(Q_{s:t},\mathcal{D}_{s})$ measures how much worse this performance is compared to the loss immediately after learning $S_{s}\sim\mathcal{D}_{s}$. Since the continual learning setting assumes that tasks are given in order, when we are given task $\mathcal{D}_{t}$ we can no longer optimize $\mathcal{L}(Q_{s},\mathcal{D}_{s})$. As such, from an optimization perspective, the best we can do at any given point is to minimize the backwards transfer loss $\mathcal{L}(Q_{s:t},\mathcal{D}_{s})$. Note that our definition of forgetting is identical to that of Lin et al. (2023). Our definition of backward transfer could be equated to either the positive part of the forgetting loss or to their definition of “generalization error”, taken only on previous tasks.
@@
The transfer loss of $Q_{s:t}$ on task $\mathcal{D}_{t}$ is defined as $\mathcal{L}(Q_{s:t},\mathcal{D}_{t})$, often referred to as the generalization loss.
Compared to the backward transfer, forward transfer (i.e., generalization) is better explored in general, and in the continual learning setting in particular, and several bounds are available (Bennani and Sugiyama, 2020; Benavides-Prado and Riddle, 2022).
Extending these notions to the full $T$ task setting, where tasks $\{1,\ldots,T\}$ appear sequentially, we define
@@
where $Q_{1:t}$ is the learned posterior based on $t$ tasks.
@@ 3 Data-dependent bounds for backward transfer
As we mentioned previously, in the continual learning setting we have no access to future tasks, and thus minimizing test forgetting reduces to minimizing the overall backward transfer as a proxy objective. In order to arrive at an upper bound on the backward transfer, we make use of concentration inequalities, relying on the change-of-measure inequality of Donsker and Varadhan (1975).
@@
$\begin{split}&\mathcal{L}(Q_{s:t},\mathcal{D}_{s})\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right].\end{split}$ (1)
The full proof of Theorem 3.1 is provided in Appendix A. The main idea is to use Lemma A.1 (the change-of-measure inequality) with $f(z)=\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))$. Unlike standard (forward) transfer results, the l.h.s. depends on the source task $s$ while the r.h.s. depends on $t$. This is necessary since we no longer have access to the training data of the source task $S_{s}$. By subtracting $\mathcal{L}(Q_{s},\mathcal{D}_{s})$ from both sides, we arrive at an upper bound on the forgetting.
Looking at the three terms in the r.h.s. of (LABEL:eq:forget-base2) we see a complex tradeoff: if ^$\hat{\mathcal{L}}(Q_{s},S_{t})$ is small, the first two terms will be small as choosing $Q_{s:t}=Q_{s}$ or a posterior with low KL-divergence will lead to low empirical loss, but the final term will be larger (we subtract a smaller number in the exponent). If, on the other hand, ^$\hat{\mathcal{L}}(Q_{s},S_{t})$ is high, we can expect the disagreement term
@@
to be small, but the first two terms will increase as achieving low ^$\hat{\mathcal{L}}(Q_{s:t},S_{t})$ requires higher KL-divergence. In both cases, if the prior $Q_{s}$ transfers well to the new task $S_{t}$, as expected for similar tasks, we see a lower r.h.s.
We note that unlike the other terms in the r.h.s. of (LABEL:eq:forget-base2), the expected loss on the previous task $\mathcal{L}(h,\mathcal{D}_{s})$ is not data-dependent. However, since this term does not depend on the current posterior $Q_{s:t}$, it has no direct effect on any optimization of the posterior.
@@
The empirical loss can be removed from the final term in (LABEL:eq:forget-base2) leading to the following result.
@@
Let $\Delta\mathcal{L}(h,s,t)\triangleq\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t})$. For any fixed $S_{s},Q_{s},\lambda_{t}>0$, the following applies uniformly for all posteriors $Q_{s:t}$ with probability at least $1-\delta$ over the choice of $S_{t}$ ($m_{t}=|S_{t}|$),
$\begin{split}\mathcal{L}(Q_{s:t},\mathcal{D}_{s})&\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\\ &+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right]\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log(1/\delta).\end{split}$ (2)
Proof of Corollary 3.2 is provided in Appendix A. The term $\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t})}\right]$ measures domain disagreement over $Q_{s}$, and can be hard to quantify in general.
For the $T$-task setting, we can bound the individual losses for each task. By using the same change of measure inequality as in Theorem 3.1, we can derive the following general bound on forgetting.
Theorem 3.3.
$\mathrm{(Forgetting)}$ For any $\lambda_{T}>0$, for any $S_{T}\sim\mathcal{D}_{T}$ and $t\in[T-1]$,
$\mathcal{L}$ $(Q_{1:T},\mathcal{D}_{t})\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})$
$+\frac{1}{\lambda_{T}}\log\mathbb{E}_{h\sim Q_{1:T-1}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{T}))}\right]$ (3)
By averaging Theorem 3.3 on all $T-1$ previous tasks, we get
Corollary 3.4.
For any fixed $Q_{1:T-1},Q_{1:T},S_{1},\ldots,S_{T}$, and $\lambda_{T}>0$,
$\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}(Q_{1:T},\mathcal{D}_{t})\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})$
$+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})$ (4)
$+\frac{1}{(T-1)\lambda_{T}}\sum_{t=1}^{T-1}\log\mathbb{E}_{Q_{1:T-1}}\left[e^{\lambda_{T}(\mathcal{L}(h,\mathcal{D}_{t})-\hat{\mathcal{L}}(h,S_{T}))}\right].$
Subtracting the loss for each task from both sides and applying the same steps as Corollary 3.2, we get
Corollary 3.5.
($\mathrm{Forgetting}$) For any fixed $\lambda_{T}>0$ and $Q_{1:T-1},S_{1},\ldots,S_{T-1}$, the following applies uniformly for all posteriors $Q_{1:T}$ with probability at least $1-\delta$ over the choice of $S_{T}$,
$F(Q_{1:T},[T])\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})-\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}(Q_{1:t},\mathcal{D}_{t})$
$+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})+\frac{\lambda_{T}K^{2}}{8m_{T}}+\frac{\log{1/\delta}}{\lambda_{T}}$
$+\frac{1}{(T-1)\lambda_{T}}\sum_{t=1}^{T-1}\log\mathbb{E}_{Q_{1:T-1}}\left[e^{\lambda_{T}\Delta\mathcal{L}(h,T,t)}\right].$ (5)
4 Oracle bounds for forward and backward transfer

### v3

@@ Abstract
In continual learning, knowledge must be preserved and re-used between tasks, maintaining good transfer to future tasks and minimizing forgetting of previously learned ones. While several practical algorithms have been devised for this setting, there have been few theoretical works aiming to quantify and bound the degree of Forgetting in general settings. For exemplar-free methods, we provide both data-dependent upper bounds that apply regardless of model and algorithm choice, and oracle bounds for Gibbs posteriors. We derive an algorithm based on our bounds and demonstrate empirically that our approach yields tight and practical bounds on forgetting for several continual learning problems and algorithms.
@@ 1 Introduction
Continual learning is a burgeoning machine learning setting where data from different tasks are presented sequentially to the learner. The usual stated goal of methods in this setting is to adapt the learner to new tasks as they appear while also preserving its performance on previous tasks (De Lange et al., 2021; Hadsell et al., 2020; Parisi et al., 2019). This performance on previous tasks is called backward transfer, or forgetting, and one of the key challenges in continual learning is avoiding Catastrophic Forgetting (Goodfellow et al., 2015; Ramasesh et al., 2020; Kirkpatrick et al., 2017), where performance on previous tasks degrades significantly as the model adapts to new tasks. Although avoiding catastrophic forgetting is desirable, much of the focus of continual learning research in recent years was on settings with a shared optimal solution. Realistically, we should not expect one model to perform optimally across all settings, for example if the data distribution changes gradually. A recent paper by Kumar et al. (2023) discusses continual learning as a computationally constrained optimization problem and argues that forgetting non-recurring information is not “catastrophic”, especially in changing environments. We will discuss this topic in our empirical evaluation.
@@
While there have been several empirical approaches in the field, there are relatively few theoretical works that explore and attempt to quantify and bound this backward transfer. Some, such as Evron et al. (2022); Lin et al. (2023); Li et al. (2024) focus on linear regression models to consider the effect of task order and similarity on forgetting. Others, such as Bennani & Sugiyama (2020); Doan et al. (2021) utilize the Neural Tangent Kernel (NTK) regime to focus on more complex task similarity measures as predictors of forgetting. Several more general works such as Benavides-Prado & Riddle (2022) apply notions of VC-dimension to arrive at more general scaling laws and data-agnostic upper bounds on forgetting for the ERM model, but may be difficult to apply for complex predictors due to the potentially large VC-dimension of models such as deep neural networks (Bartlett et al., 2019). We note that many of the known results (Bennani & Sugiyama, 2020; Evron et al., 2022; Benavides-Prado & Riddle, 2022) provide upper bound on forgetting for the training data, though there are previous results on forgetting for test data as well (Doan et al., 2021; Lin et al., 2023; Li et al., 2024) for specific models. In particular, Lin et al. (2023) provide an exact characterization of forgetting for the Stochastic Gradient Descent (SGD) algorithm under specific assumptions in the linear regression setting. Bennani & Sugiyama (2020); Doan et al. (2021) derive bounds for more general settings for very wide architectures that follow the NTK regime, providing bounds specifically for the SGD algorithm as well as a continual learning variant known as Orthogonal Gradient Descent (OGD).
@@
In this work, we explore upper bounds on forgetting that apply to general architectures and optimization algorithms . We note that existing results are restricted to specific setups (e.g.linear regression) and algorithms (e.g. SGD). We use the PAC-Bayes (McAllester, 1999; Catoni, 2004; Alquier, 2024) framework to derive and analyze upper bounds on backward transfer, focusing on the Gibbs posterior (Casella & George, 1992). In Section 3 we derive general bounds for the two-task and multiple-task settings without making any model assumptions. We then focus our discussion in Section 4 on oracle bounds for the Gibbs posterior in general, and under specific assumptions of task similarity. We then proceed to empirically demonstrate our bounds [^1] Code is available at https://github.com/lioritan/continual_forgetting_pb for both a Variational Inference (VI) (Hoffman et al., 2013) based algorithm as well as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), on several continual learning problems in Section 5. To the best of our knowledge, our bounds are the first general PAC-Bayes upper bounds on forgetting in continual learning that are applicable regardless of model choice and task distribution, as well as the first established forgetting upper bounds for the EWC algorithm. Table 1 provides an overview of existing settings and limitations for known results, as well as our own results.
Table 1: Settings covered by existing test forgetting bounds as well as our bounds (final two rows). Methods marked with * use added noise during bound estimation for deterministic architectures.
Algorithm Architecture Task construction Reference
SGD Linear, overparametrized Linear regression Lin et al. (2023)
SGD Linear, Mixture-of-experts Linear regression Li et al. (2024)
SGD and OGD NTK regime Any Bennani & Sugiyama (2020)
SGD and OGD NTK regime Any Doan et al. (2021)
Data-Agnostic Any $\mathcal{F}$-related tasks Benavides-Prado & Riddle (2022)
Any Any* Any Corollary 3.2
Emprirical Gibbs Any* Any Theorems 4.3,4.4
@@ 2 Problem definition
Figure 1: Continual Learning problem. A data-free prior $P$ is adapted via an empirical sample $S_{1}\sim\mathcal{D}_{1}$ to a posterior $Q_{1}$. This posterior serves as the new prior for the next task and so on until we reach a final posterior $Q_{1:T}$.
@@
The expected loss of a given hypothesis $h\in\mathcal{H}$ is defined as $\mathcal{L}(h,\mathcal{D})\triangleq\mathbb{E}_{z\in\mathcal{D}}\ell(h,z)$. The empirical loss of a hypothesis w.r.t. a sample $S$ is defined as $\hat{\mathcal{L}}(h,S)\triangleq\frac{1}{m}\sum_{j=1}^{m}\ell(h,z_{j})$.
In the following Section, and in Section 4, we first consider only two tasks from distributions $\mathcal{D}_{s}$ and $\mathcal{D}_{t}$, referring to source and target. Within the classic PAC-Bayes setting (McAllester, 1999; Alquier, 2024), let $Q_{s}$ be a distribution over the set of hypotheses learned by some algorithm $J_{s}$ over $S_{s}\sim\mathcal{D}_{s}$ and a data-free prior hypothesis distribution $P$, such that $Q_{s}=J_{s}(S_{s},P).$ We then proceed to utilize another algorithm $J_{t}$ that operates on $S_{t}\sim\mathcal{D}_{t}$, making use of $Q_{s}$, such that $Q_{s:t}=J_{t}(S_{t},Q_{s}).$ For now, we make no assumptions on $J_{s}$ and $S_{t}$ other than their inputs. Of particular note is that algorithm $J_{t}$ is exemplar-free (i.e. it has no access to data from $\mathcal{D}_{s}$), and that task $\mathcal{D}_{t}$ is chosen independently of the prior $Q_{s}$, meaning that task choice is independent of the learning process.
@@
$F(Q_{s:t},\mathcal{D}_{s})$ $\triangleq\mathrm{BWT}(Q_{s:t},\mathcal{D}_{s})-\mathbb{E}_{h\sim Q_{s}}\left[\mathcal{L}(h,\mathcal{D}_{s})\right]=\mathcal{L}(Q_{s:t},\mathcal{D}_{s})-\mathcal{L}(Q_{s},\mathcal{D}_{s}).$
Intuitively, the backward transfer $\mathcal{L}(Q_{s:t},\mathcal{D}_{s})$ measures the performance of the updated model on the previous task $\mathcal{D}_{s}$, and the forgetting $F(Q_{s:t},\mathcal{D}_{s})$ measures how much worse this performance is compared to the loss immediately after learning $S_{s}\sim\mathcal{D}_{s}$. Since the continual learning setting assumes that tasks are given in order, when we are given task $\mathcal{D}_{t}$ we can no longer optimize $\mathcal{L}(Q_{s},\mathcal{D}_{s})$. As such, from an optimization perspective, the best we can do at any given point is to minimize the backward transfer loss $\mathcal{L}(Q_{s:t},\mathcal{D}_{s})$. Our definition of forgetting is identical to that of Lin et al. (2023) (but for non-deterministic hypothesis). Our definition of backward transfer could be equated to either the positive part of the forgetting loss or to their definition of “generalization error”, taken only on previous tasks.
@@
The transfer loss of $Q_{s:t}$ on task $\mathcal{D}_{t}$ is defined as $\mathcal{L}(Q_{s:t},\mathcal{D}_{t})$, also known as the generalization loss.
Compared to the backward transfer, forward transfer (i.e., generalization) is better explored in general, and in the continual learning setting in particular, and several bounds are available (Bennani & Sugiyama, 2020; Benavides-Prado & Riddle, 2022).
We also note that several existing PAC-Bayes bounds for generalization can be extended or altered for forward transfer in the continual setting: Haddouche & Guedj (2022) derive generalization bounds for online learning, and both Haddouche & Guedj (2023) and Chugg et al. (2023) derive anytime bounds that can be modified for generalization in the continual learning setting. While not directly relevant, we also note that there are existing PAC-Bayes bounds for generalization in the context of meta-learning (Pentina & Lampert, 2014; 2015; Amit & Meir, 2018), that can be considered as an offline version of the continual learning setting. To the best of our knowledge, there are no bounds on forgetting in the context of meta-learning.
Extending these notions to multiple task setting, where tasks $\{1,\ldots,T\}$ appear sequentially, we define
@@
where $Q_{1:t}$ is the learned posterior based on $t$ tasks. See also Appendix A for a detailed summary of notation.
@@ 3 Data-dependent bounds for backward transfer
As we mentioned previously, in the continual learning setting we have no access to future tasks, and thus minimizing test forgetting reduces to minimizing the overall backward transfer as a proxy objective. Since any upper bound on backward transfer can be converted to an upper bound on forgetting, we focus our efforts on upper bounds for backward transfer. In order to arrive at an upper bound on the backward transfer, we make use of concentration inequalities, relying on the change-of-measure inequality of Donsker & Varadhan (1975).
@@
$\begin{split}&\mathcal{L}(Q_{s:t},\mathcal{D}_{s})\leq\hat{\mathcal{L}}(Q_{s:t},S_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}(\mathcal{L}(h,\mathcal{D}_{s})-\hat{\mathcal{L}}(h,S_{t}))}\right].\end{split}$ (1)
The full proof of Theorem 3.1 is provided in Appendix B. The proof follows the same basic techniques and structure of PAC-Bayes generalization upper bounds (McAllester, 1999; Alquier, 2024). The main idea is to use Lemma B.1 (the change-of-measure inequality) with $f(z)=\lambda_{t}(\mathcal{L}(z,\mathcal{D}_{s})-\hat{\mathcal{L}}(z,S_{t}))$. Unlike standard (forward) transfer results, the l.h.s. depends on the source task $s$ while the r.h.s. depends on task $t$. This is necessary since we no longer have access to the training data of the source task $S_{s}$. By subtracting $\mathcal{L}(Q_{s},\mathcal{D}_{s})$ from both sides, we arrive at an upper bound on the forgetting.
Looking at the three terms in the r.h.s. of (LABEL:eq:forget-base2) we see a complex tradeoff: if $\hat{\mathcal{L}}(Q_{s},S_{t})$ is small, the first two terms will be small as choosing $Q_{s:t}=Q_{s}$ or a posterior with low KL-divergence will lead to low empirical loss, but the final term will be larger (we subtract a smaller number in the exponent). If, on the other hand, $\hat{\mathcal{L}}(Q_{s},S_{t})$ is high, we can expect the disagreement term
@@
to be small, but the first two terms will increase as achieving low $\hat{\mathcal{L}}(Q_{s:t},S_{t})$ requires higher KL-divergence. In both cases, if the prior $Q_{s}$ transfers well to the new task $S_{t}$, as expected for similar tasks, we see a lower r.h.s.
We note that unlike the other terms in the r.h.s. of (LABEL:eq:forget-base2), the expected loss on the previous task $\mathcal{L}(h,\mathcal{D}_{s})$ is not data-dependent. However, since the disagreement term does not depend on the current posterior $Q_{s:t}$, it has no direct effect on any optimization of the posterior.
@@
For the $T$-task setting, we can bound the individual losses for each task. By using the same change of measure inequality as in Theorem 3.1, and averaging over previous tasks, we get
@@
($\mathrm{Forgetting}$) Let
$\Delta\mathcal{L}(h,s,t)\triangleq\mathcal{L}(h,\mathcal{D}_{s})-\mathcal{L}(h,\mathcal{D}_{t}).$
For any fixed $\lambda_{T}>0$ and $Q_{1:T-1},S_{1},\ldots,S_{T-1}$, the following applies uniformly for all posteriors $Q_{1:T}$ with probability at least $1-\delta$ over the choice of $S_{T}$ ($m_{t}=|S_{t}|$),
$F(Q_{1:T},[T])$ $\leq\hat{\mathcal{L}}(Q_{1:T},S_{T})-\frac{1}{T-1}\sum_{t=1}^{T-1}\mathcal{L}(Q_{1:t},\mathcal{D}_{t})+\frac{1}{\lambda_{T}}D_{\mathrm{KL}}(Q_{1:T}||Q_{1:T-1})$
$+\frac{\lambda_{T}K^{2}}{8m_{T}}+\frac{\log{1/\delta}}{\lambda_{T}}+\frac{1}{(T-1)\lambda_{T}}\sum_{t=1}^{T-1}\log\mathbb{E}_{Q_{1:T-1}}\left[e^{\lambda_{T}\Delta\mathcal{L}(h,t,T)}\right].$ (2)
The proof of Corollary 3.2 appears in Appendix B.
4 Oracle bounds for backward transfer